1.  The following performance report was prepared for Dale Manufacturing for the month of April.

                                     Actual Results Static Budget Variance

Sales units                   100,000           80,000             20,000F
Sales dollars                $190,000         $160,000         $30,000F
Variable costs             125,000           96,000             29,000U
Fixed costs                  45,000             40,000             5,000U
Operating income       $ 20,000          $ 24,000          $ 4,000U

Using a flexible budget, Dale’s total sales-volume variance is:

a. $6,000 F b. $16,000 F c. $4,000 U d. $20,000 U

2. MinnOil performs oil changes and other minor maintenance services (e.g., tire pressure checks) for cars. The company advertises that all services are completed within 15 minutes for each service. On a recent Saturday, 160 cars were serviced resulting in the following labor variances: rate, $19 unfavorable; efficiency, $14 favorable. If MinnOil’s standard labor rate is $7 per hour, determine the actual wage rate per hour and the actual hours worked. Wage Rate Hours Worked.

a. $7.45---42.00 b. $7.50---38.00 c. $6.55---42.00 d. $6.67---42.71

3. Frisco Company recently purchased 108,000 units of raw material for $583,200. Three units of raw materials are budgeted for use in each finished good manufactured, with the raw material standard set at $16.50 for each completed product. Frisco manufactured 32,700 finished units during the period just ended and used 99,200 units of raw material. If management is concerned about the timely reporting of variances in an effort to improve cost control and bottom-line performance, the materials purchase price variance should be reported as:

a. $9,920 F b. $6,050 U c. $10,800 F d. $10,800 U
Answer #1

1. Sales volume variance :

= (Budgeted quantity - Actual quantity) × Budgeted profit(W1)

=(80000 - 100000) ×0.3 = 6000 F

So option A is right.

W1:

Budget profit;

Sales 160000
Variable cost 96000
Fixed cost 40000
Profit 26000
Profit per unit 26000/80000 =0.3


2. Actual hours worked and Actual Rate :

Calculating Actual hours,

Labour efficiency variance =

(Standard hours - Actual hours)×Standard Rate = 14(given)

=[(160×15/60) - Actual hours) ×7 =14

=[40 - Ac​​tual hours] =14÷7

Actual hours = 40 - 2 = 38 hours.

Calculating Actual Rate,

Labour Rate varience =

(Standard Rate - Actual Rate) × Actual Hours = - 19(given)

=( 7 - Actual Rate) × 38(calculated) = - 19

= 7 - Actual Rate = - 19 ÷ 38

Actual Rate = 7 + 19/38 = $ 7.5.

​​​​​​Therefore, option B is correct.

3. Material price varience =

(Standard price - Actual price) *Actual quantity

=[(16.5/3) - (583200/108000)] × 99200

=(5.5 - 5.4) ×99200

= 9920 Favourable.

So option A is correct.